# Simulating Online Social Media Conversations on Controversial Topics Using AI Agents Calibrated on Real-World Data

### v1

@@ 3 Experimental Design
  • ReverseChronoFollowersPopularity: the default setting, where recent posts from followed users are shown, ranked by popularity, with some exposure to external users.
  • ContentRecSys: a random baseline, where posts are sampled randomly from all platform content.
Opinion modelling
In addition to the simulation-driven opinion evolution, we are considering a well-grounded in literature approach: the Friedkin–Johnsen model (Friedkin and Johnsen 1990) is a foundational framework in opinion dynamics that extends the classical DeGroot(DeGroot 1974b) model by incorporating the concept of individual resistance to change. In this model, each agent updates their opinion as a weighted combination of their neighbors’ views and their own initial belief, with a susceptibility parameter determining the balance between external influence and internal stubbornness.
Coalition distribution in the population
At initialization, users were assigned political leanings by sampling from real-world data, reproducing the coalition distribution observed in the source dataset. As shown in Figure 1, the population is imbalanced: the Right coalition dominates, Centre-Left and Third Pole are of comparable size, and M5S is smaller with low variability across simulations.
Toxicity analysis
4 Results
4.1 In-group and out-group interactions between agents
We begin by examining patterns of in-group and out-group interactions among agents—specifically, the frequency with which agents engage with others affiliated with the same political coalition versus those aligned with opposing coalitions. To assess the similarity between simulated and real-world behavior, we aggregate interactions from 10 independent runs for each parameter configuration (model, network initialization, and recommender system) into a single dataset. We then compute the proportion of in-group and out-group interactions for each coalition, applying the same procedure to the empirical interaction data used to initialize the simulations. In particular, we focus on replies. Finally, we evaluate correspondence by calculating Pearson correlations: four values capturing in-group alignment (one per coalition) and twelve values capturing out-group interactions (three per coalition).
Figure 2 reports the distribution of correlations between simulated and real in-group interaction frequencies, disaggregated by model type and network initialization. Focusing first on the uncensored model (top row), we observe that the recommender system exerts the strongest influence. In particular, the Reverse Chronological system substantially improves alignment with the real data, raising the median correlation to 0.85, an increase of almost 30 percentage points over the Random system (median = 0.6). While some runs with the Random system achieve similar levels of accuracy, its outcomes are far less stable, with correlations falling as low as –0.28 when simulations begin from an empty network. This gap narrows under the fully connected initialization, where the two recommenders yield similar median correlations; however, variance remains consistently higher in the Random condition, except for Llama2-70B in the Reverse Chronological system that exhibits outliers with negative correlation. Overall, however, the simulations tend to reproduce fairly well the extent to which users interact homophilously, at least on average. To illustrate the analysis in more detail, Figure 3 presents a case study of a single experimental configuration (Llama3.2-3B, empty initialization, Random recommender). Here, we compare the distribution of simulated in-group interactions for each coalition against the corresponding empirical proportions. The summary results in Figure 2 are then derived by computing, for each run, the Pearson correlation between the four coalition-level values produced by the simulation and those observed in the real-world dataset.
Figure 2: Distribution of the correlation of in-group interactions between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation.
Figure 5 illustrates these dynamics through a detailed example. For a given configuration (Llama3.2-3B, empty network, Random recommender), we compare the distribution of simulated out-group interactions for each coalition pair with those observed in the real data. The left panel displays the percentage of simulated interactions flowing from one coalition to another, while the right panel presents the corresponding heatmap derived from the empirical dataset. The summary correlations reported in Figure 4 are then obtained by comparing, for each run, the twelve simulated values (off-diagonal entries of the interaction matrix) with their twelve real-world counterparts.
Figure 4: Distribution of the correlation of out-group interactions between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation.
Figure 5: Example of comparison between simulated (10 runs across 1 configuration) and real data. On the left, the heatmap of out-group interactions percentages for each coalition pair (using Llama3.2-3B, empty initial network, and random recommender system). On the right, the same heatmap in the real-world dataset.
4.2 In- and out-group toxic behaviour
We next turn to the analysis of comment toxicity in the simulations. Our goal is to evaluate whether the simulated environments reproduce the patterns of hostile communication observed in the real-world data. To this end, we measure toxicity by computing the toxicity score of all replies and extracting the $95^{\mathsf{th}}$ percentile for each parameter configuration. We then distinguish between toxicity directed at members of the same coalition (in-group toxicity) and toxicity directed at members of other coalitions (out-group toxicity).
To assess how faithfully the simulations capture these patterns, we adopt the same strategy used in the interaction analyses: we correlate the simulated toxicity values with those observed in the empirical dataset used to initialize the runs. Figure 6 reports the resulting correlations for in-group toxicity, broken down by recommender system and network initialization strategy.
Across most configurations, we observe high variability in the simulations. The only clear exception is Llama2-70B with a fully connected initialization, which yields a compact distribution of correlations with relatively high values under both recommendation strategies. In other settings, the most stable configuration is the combination of the ReverseChronological recommender with an empty starting network, which achieves the highest median correlation overall.
Turning to out-group toxicity, we find less variability across runs but also a reversed trend with respect to network initialization. With an empty network initialization, the ReverseChronological algorithm produces lower correlations with the real data, whereas under the fully connected initialization it performs better. The choice of model appears to have little influence on the outcome, as median correlation values remain largely comparable across the two uncensored models.
Taken together, these findings suggest that simulations struggle to replicate patterns of toxic behavior among online users, despite being based on uncensored models. As in the case of interaction analyses, illustrative examples of preliminary investigations are presented in Figures 7 and 9.
Figure 6: Distribution of the correlation of in-group toxicity between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation.
Figure 8: Distribution of the correlation of inter-group toxicity between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation.
Figure 9: Example of comparison between simulated (10 runs across 1 configuration) and real data. On the left, the heatmap of out-group toxicity percentages for each coalition pair (using Llama3.2-3B, empty initial network, and random recommender system). On the right, the same heatmap in the real-world dataset.
4.3 Opinion dynamics
As shown in Appendix D, these dynamics are consistent across topics: opinions progressively converge toward neutral values, reflecting a decline in polarization over time. Whether this trend persists or stabilizes with longer simulations remains an open question. Importantly, the same pattern is observed across models, network structures, and recommender systems, underscoring the robustness of the result.
5 Discussion and Conclusion
LLMs have emerged as a promising tool for simulating agents in virtual environments. The main goal of this work was to analyze the behavior of LLM-based agents in the context of online social media platforms. To this end, we extended the Y simulator by integrating mechanisms for opinion evolution and introducing a realistic user initialization procedure grounded in empirical data from the 2022 Italian political context. This framework enabled us to systematically test how different generative models, network initialization strategies, and recommender systems affect the fidelity of the simulations when compared with real-world observations.
At the same time, several limitations should be acknowledged. The 21 simulated days were sufficient for a network structure to begin forming, but too short to capture longer-term or emergent dynamics. For instance, actions such as unfollow were almost absent, and the full effects of recommendation algorithms may not have materialized given the relatively unstructured networks in the early phases of simulation. With respect to opinion evolution, LLM-assigned scores were broadly consistent with those produced by traditional models, both showing a tendency to converge toward neutral positions. However, the relatively short duration leaves open whether these trends would stabilize, polarize, or diverge over longer time horizons. Our choice of simulation length was ultimately constrained by available computational resources.
Future developments should focus on enriching agent personalization. Incorporating elements such as emotional reasoning, susceptibility to influence, or varying levels of trust in consumed information could enable more realistic opinion and interaction dynamics. Another important extension is the inclusion of external shocks—such as political crises, scandals, or major public statements—to evaluate how agents respond to events that typically shape online discourse. More systematic comparisons with empirical data are also necessary to better evaluate the realism of emergent behaviors. Finally, expanding beyond the Italian political context would allow us to assess the generalizability of the approach and test whether the observed dynamics hold across different sociopolitical settings.
In conclusion, integrating LLMs as agents in social simulations represents an important step toward more realistic modeling of online environments, particularly with respect to language, interactions, and content generation. However, replicating more heterogeneous phenomena, such as the spread of misinformation, will require further advances in behavioral modeling. Our study contributes to the growing exploration of LLM-driven simulations, providing a foundation for investigating complex social dynamics under controlled conditions.
6 Code and Data Availability
All code and prompts to replicate the analyses will be made available on GitHub.
Appendix
Appendix A Computational resources
[…]
Appendix B Prompts
@@ C.4 Third Pole
Appendix D Additional Plots
D.1 Opinion Shifts

### v2

@@ Introduction
  • ReverseChronoFollowersPopularity: the default setting (henceforth Default), where recent posts from followed users are shown, ranked by popularity, with some exposure to external users.
  • ContentRecSys: a random baseline (henceforth Random), where posts are sampled randomly from all platform content.
Opinion modelling
In addition to the simulation-driven opinion evolution, we are considering a well-grounded in literature approach: the Friedkin–Johnsen model (Friedkin and Johnsen 1990) is a foundational framework in opinion dynamics that extends the classical DeGroot (DeGroot 1974b) model by incorporating the concept of individual resistance to change. In this model, each agent updates their opinion as a weighted combination of their neighbors’ views and their own initial belief, with a susceptibility parameter determining the balance between external influence and internal stubbornness. We choose the Friedkin–Johnsen model because it generalizes the DeGroot framework while relaxing some of its restrictive assumptions, allowing opinions to stabilize without necessarily converging top consensus. Moreover, unlike bounded-confidence models such as Hegselmann–Krause (Hegselmann and Krause 2002), it does not impose explicit clustering, making it well suited for scenarios where influence is continuous and heterogeneous rather than strictly segmented.
Coalition distribution in the population
At initialization, users were assigned political leanings by sampling from real-world data, reproducing the coalition distribution observed in the source dataset. As shown in Figure 1, the population is imbalanced: the Right coalition dominates, Centre-Left and Third Pole are of comparable size, and M5S (Movimento 5 Stelle) is smaller with low variability across simulations. More information about the Italian political coalitions can be found in Appendix A.
Toxicity analysis
In-group and out-group interactions between agents
We begin by examining patterns of inter-group interactions among agents, divided into in-group and out-group interactions; specifically, the frequency with which agents engage with others affiliated with the same political coalition versus those aligned with opposing coalitions. To assess the similarity between simulated and real-world behavior, we aggregate interactions from 10 independent runs for each parameter configuration (model, network initialization, and recommender system) into a single dataset. We then compute the proportion of in-group and out-group interactions for each coalition, applying the same procedure to the empirical interaction data used to initialize the simulations. In particular, we focus on replies. Finally, we evaluate correspondence by calculating Pearson correlations: four values capturing in-group alignment (one per coalition) and twelve values capturing out-group interactions (three per coalition). For in-group interactions, we compute the correlations between the diagonals of the simulated and empirical interaction matrices. These values capture the degree to which the simulated model reproduces the proportion of interactions occurring among users supporting the same political coalition.
For out-group interactions, we instead focus on the off-diagonal elements of the interaction matrices, which represent interactions between users of different political leanings. In this case, the correlations quantify how well the simulated percentages of cross-coalition interactions align with those observed in real data.
Figure 2 reports the distribution of correlations between simulated and real in-group interaction frequencies, disaggregated by model type and network initialization. Focusing first on the smaller model (top row), we observe that the recommender system exerts the strongest influence. In particular, the Default system substantially improves alignment with the real data, raising the median correlation to 0.85, an increase of almost 30 percentage points over the Random system (median = 0.6). While some runs with the Random system achieve similar levels of accuracy, its outcomes are far less stable, with correlations falling as low as –0.28 when simulations begin from an empty network. This gap narrows under the fully connected initialization, where the two recommenders yield similar median correlations; however, variance remains consistently higher in the Random condition, except for Llama2-70B in the Default system that exhibits outliers with negative correlation. Overall, however, the simulations tend to reproduce fairly well the extent to which users interact homophilously, at least on average. To illustrate the analysis in more detail, Figure 3 presents a case study of a single experimental configuration (Llama3.2-3B, empty initialization, Random recommender). Here, we compare the distribution of simulated in-group interactions for each coalition against the corresponding empirical proportions. The summary results in Figure 2 are then derived by computing, for each run, the Pearson correlation between the four coalition-level values produced by the simulation and those observed in the real-world dataset.
Figure 2: Distribution of the Pearson correlations of in-group interactions between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation. Overall, approximately 91% of the correlations are not significant ($p>0.05$). Similar results are obtained using Spearman $\rho$.
Figure 4: Distribution of the Pearson correlations of out-group interactions between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation. Overall, approximately 66% of the correlations are not significant ($p>0.05$). Similar results are obtained using Spearman $\rho$.
To clarify how out-group interaction patterns are compared between simulations and real data, we briefly outline the procedure here and provide a concrete example in Appendix A (Figure 10). For a fixed configuration (Llama3.2-3B, empty initial network, Random recommender), we construct an inter-coalition interaction matrix from the simulations and an analogous matrix from the empirical dataset, both normalized to represent the proportion of interactions flowing from each coalition to every other coalition. For each simulation run, we extract the twelve off-diagonal entries of the simulated matrix, corresponding to out-group interactions, and directly compare them with the matching twelve empirical values. This produces a single correlation score per run, quantifying how closely the simulated interaction structure aligns with the observed one. The summary correlations obtained through this procedure, aggregated across runs and configurations, are reported in Figure 4.
In- and out-group toxic behaviour
We now explore the toxicity of the comments in the simulations. Our goal is to evaluate whether the simulated environments produce similar patterns of hostile communication observed in the real-world data. To this end, we measure toxicity by computing the toxicity score of all replies and extracting the $95^{\mathsf{th}}$ percentile for each parameter configuration. We then distinguish between toxicity directed at members of the same coalition (in-group toxicity) and toxicity directed at members of other coalitions (out-group toxicity).
To assess how faithfully the simulations capture these patterns, we adopt the same strategy used in the interaction analyses: we correlate the simulated toxicity values with those observed in the empirical dataset used to initialize the runs. Figure 5 reports the resulting correlations for in-group toxicity, divided by recommender system and network initialization strategy.
Across most configurations, we observe high variability in the simulations. The only clear exception is Llama2-70B with a fully connected initialization, which yields a compact distribution of correlations with relatively high values under both recommendation strategies. In other settings, the most stable configuration is the combination of the Default recommender with an empty starting network, which achieves the highest median correlation overall.
Turning to out-group toxicity (Figure 6), we find less variability across runs and also a similar trend with respect to network initialization compared to in-group analysis. With an empty network initialization, the Default algorithm produces high correlations with the real data, whereas under the fully connected initialization it performs worse. The choice of model appears to have little influence on the outcome, as median correlation values remain largely comparable across the two uncensored models. Taken together, these findings suggest that simulations struggle to replicate patterns of toxic behavior among online users, despite being based on uncensored models. As in the case of interaction analyses, illustrative examples of preliminary investigations are presented in Figure 7 and in the Appendix A (Figure 11). We see that, while the real distribution of in-group toxicity is evenly present across all political coalitions, the models fail to reproduce this phenomenon, with the most significant cases being M5S coalition and Center-Left. Similarly, the same pattern emerges for overall out-group toxicity, where some trends are correctly replicated, while others (e.g., Third Pole vs. M5S) are greatly over estimated or under estimated (the visual presentation of this is visible in Appendix A, Figure 11).
Figure 5: Distribution of the Pearson correlations of in-group toxicity between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation. Overall, approximately 98% of the correlations are not significant ($p>0.05$). Similar results are obtained using Spearman $\rho$.
Figure 6: Distribution of the Pearson correlations of out-group toxicity between the simulation and the real data, divided by model and network initialization strategy. Each dot is a simulation. Overall, approximately 97% of the correlations are not significant ($p>0.05$). Similar results are obtained using Spearman $\rho$.
Opinion dynamics
As shown in Appendix A, these dynamics are consistent across topics: opinions progressively converge toward neutral values, reflecting a decline in polarization over time. Whether this trend persists or stabilizes with longer simulations remains an open question. Importantly, the same pattern is observed across models, network structures, and recommender systems, underscoring the robustness of the result.
Discussion and Conclusion
LLMs have emerged as a promising tool for simulating agents in virtual environments. The main goal of this work was to provide an exploratory analysis of the behavior of LLM-based agents in the context of online social media platforms. To this end, we extended the Y simulator by integrating mechanisms for opinion evolution and introducing a realistic user initialization procedure grounded in empirical data from the 2022 Italian political context. This framework enabled us to systematically test how different generative models, network initialization strategies, and recommender systems affect the fidelity of the simulations when compared with real-world observations.
At the same time, several limitations should be acknowledged. The 21 simulated days were sufficient for a network structure to begin forming, but too short to capture longer-term or emergent dynamics. For instance, actions such as unfollow were almost absent, and the full effects of recommendation algorithms may not have materialized given the relatively unstructured networks in the early phases of simulation. Future work should address this by implementing more faithful initialization techniques, such as Stochastic Block Models, to generate networks with a realistic density and predefined community structure. This will enable a more robust and nuanced study of emergent dynamics within established social environments. With respect to opinion evolution, LLM-assigned scores were broadly consistent with those produced by traditional models, both showing a tendency to converge toward neutral positions. However, the relatively short duration leaves open whether these trends would stabilize, polarize, or diverge over longer time horizons. Our choice of simulation length was ultimately constrained by available computational resources.
Although simulated interaction and toxicity patterns were not significantly correlated with empirical data, this result should be interpreted in light of the exploratory and short-horizon nature of the simulations rather than as a failure of the modeling approach. The lack of significant correlations is in contrast with previous findings, particularly Gao et al. (2023), showing that it is crucial for both traditional opinion dynamics models and LLM-based agents to have longer time horizons and more structured networks to produce stable, macro-level alignment with empirical observations. From this perspective, our results suggest that current LLM-based simulations can be considered still in their early stage of development and far from being suited for direct quantitative prediction of real-world interaction patterns.
Nevertheless, despite the limited temporal scope of our simulations, our results are consistent with those reported by Chuang et al. (2024) with respect to opinion evolution. In particular, agents exhibit a clear tendency to converge toward neutral positions. This behavior aligns with previous findings showing that, in the absence of explicitly modeled cognitive biases, opinion dynamics tend to stabilize rather than polarize.
Future developments should focus on enriching agent personalization and studying the robustness of these systems. Incorporating elements such as emotional reasoning, susceptibility to influence, or varying levels of trust in consumed information could enable more realistic opinion and interaction dynamics. We also relied on a fixed prompt structure. Considering the critical role that prompt sensitivity plays in LLM–based experiments, future work should perform ablation studies to assess how sensitive agent behavior is to variations in prompt phrasing. Another important extension is the inclusion of external shocks—such as social crises (Di Giovanni et al. 2022), scandals, or major public statements—to evaluate how agents respond to events that typically shape online discourse. It could also be of interest to see the impact of simulated harmful content, such as conspiracy theories (Corso et al. 2025b), misinformation, disinformation (Nogara et al. 2026), or political advertisement (Pierri 2023). More systematic comparisons with empirical data are also necessary to better evaluate the realism of emergent behaviors. Finally, expanding beyond the Italian political context would allow us to assess the generalizability of the approach and test whether the observed dynamics hold across different sociopolitical settings. In conclusion, integrating LLMs as agents in social simulations represents an important step toward more realistic modeling of online environments, particularly with respect to language, interactions, and content generation. However, replicating more heterogeneous phenomena, such as the spread of misinformation, will require further advances in behavioral modeling. Our study contributes to the growing exploration of LLM-driven simulations, providing a foundation for investigating complex social dynamics under controlled conditions.
Ethical Statement
The deployment of Large Language Models as AI social agents raises numerous ethical considerations that are currently the subject of intense scrutiny by the interdisciplinary research community. The extraordinary capabilities of these models to generate text have led several scientists to envision alarming scenarios in which the seamless integration of AI agents into the online social discourse may facilitate the dissemination of harmful content, the spread of misinformation, and the propagation of ‘semantic garbage’, ultimately damaging our societies (Floridi and Chiriatti 2020; Weidinger et al. 2022; Hendrycks et al. 2023). As a result, any research exploring the characteristics of LLMs as social agents could, directly or indirectly, contribute knowledge that might be exploited to implement and deploy LLM-based technologies for malicious purposes. While recognizing this risk, we also believe that conducting research on LLM-based agents is essential to assess potential risks and to guide efforts aimed at developing strategies to mitigate them. Our study contributes positively to deepen our understanding of how LLMs react to social stimuli.
Even when deploying LLM-based agents for ethical purposes, trade-offs between the obtained benefit and the high level of power consumption required to run them should be carefully considered (Bender et al. 2021).
Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy). Francesco Pierri is partially supported by PNRR-PE-AI FAIR project funded by the NextGeneration EU program.
Appendix A Appendix
Computational Resources
[…]
Code and Data Availability
All code and prompts to replicate the analyses are available on GitHub: github.com/elisacomposta/YAnalysis.
@@ Appendix A Appendix
Dataset Information
The ITA-ELECTION-2022 is a multi platform social media dataset described in Pierri et al. (2023). For the purpose of our study we employed only the Twitter portion of the dataset composed of 19,087,594 tweets posted by 618,089 unique accounts. The data was collected with the Twitter API over the period of the General Italian Elections, from September to October 2022. Political affiliation was inferred following the methodology proposed in (Pierri 2024). The approach relies on a set of 471 Twitter accounts belonging to elected members of the Italian Senate and Chamber of Deputies, grouped into four major political coalitions and provided as part of the ITA-ELECTION-2022 dataset. These accounts are used as labeled reference points to propagate political labels to ordinary users. Specifically, users who retweeted at least one politician were assigned a political affiliation based on the coalition they retweeted most frequently. User-level toxicity was derived by aggregating the toxicity scores of tweets authored by each user.
Prompts
@@
Additional Plots
Inter-Group Interactions Example
Figure extending the analysis presented in Section In-group and out-group interactions between agents
Figure 10: Example of comparison between simulated (10 runs across 1 configuration) and real data. On the left, the heatmap of inter-group interactions percentages for each coalition pair (using Llama3.2-3B, empty initial network, and random recommender system). On the right, the same heatmap in the real-world dataset.
Inter-Group Toxicity Example
Figure extending the analysis presented in Section In- and out-group toxic behaviour
Figure 11: Example of comparison between simulated (10 runs across 1 configuration) and real data. On the left, the heatmap of inter-group toxicity percentages for each coalition pair (using Llama3.2-3B, empty initial network, and random recommender system). On the right, the same heatmap in the real-world dataset.
Opinion Shifts
Figures extending the analysis presented in Section Opinion dynamics